# Benchmarking Pre-trained Large Language Models’ Potential Across Urdu NLP tasks

### v1

@@ Abstract
Large Language Models (LLMs) pre-trained on multilingual data have revolutionized natural language processing research, by transitioning from languages and task specific model pipelines to a single model adapted on a variety of tasks. However majority of existing multilingual NLP benchmarks for LLMs provide evaluation data in only few languages with little linguistic diversity. In addition these benchmarks lack quality assessment against the respective state-of the art models. This study presents an in-depth examination of prominent LLMs; GPT-3.5-turbo, Llama2-7B-Chat, Bloomz 7B1 and Bloomz 3B, across 14 tasks using 15 Urdu datasets, in a zero-shot setting, and their performance against state-of-the-art (SOTA) models, has been compared and analysed. Our experiments show that SOTA models surpass all the encoder-decoder pre-trained language models in all Urdu NLP tasks with zero-shot learning. Our results further show that LLMs with fewer parameters, but more language specific data in the base model perform better than larger computational models, but low language data.
@@ 1 Introduction
Our study, focuses on evaluating the potential of both closed and open LLMs for supporting Urdu, a low resource language with limited data coverage in LLM’s pre-training. In our experiments we utilize GPT3.5 turbo by OpenAI, Llama 2 by Meta and Bloomz 3B and 7B1 by Big Science in zero-shot setting, and perform evaluation on 14 Urdu NLP tasks analysing their performances with the existing SOTA models. To the best of our knowledge, this is the first in depth evaluation of prominent LLMs in Urdu Language context.
@@
Cyber bullying Identification Cyberbullying corpus (Adeeba et al., 2024) 12,759 2480
@@
Machine Translation(Quran) English-Urdu Religious Parallel Corpus (Jawaid and Zeman, 2011) 6,414 200
Machine Translation(Bible) English-Urdu Religious Parallel Corpus (Jawaid and Zeman, 2011) 7,957 257
@@ 2 Approach
For benchmarking of Urdu NLP tasks, we perform experiments using GPT 3.5, Bloomz 3B and 7B1 and Llama 2 in zero-shot setting and comparatively analyse the results with the respective SOTA models. Model selection was based on factors like accessibility (open/closed), infrastructure requirement, performance, language support. GPT 3.5 was selected because of its superior performance on English tasks. Among open models, popular multilingual models i.e. Llama 2 and Bloomz were evaluated. Due to budget limitations and lack of Urdu data in the pre-training, other closed models were not investigated.
The evaluation of LLMs involved prompting and significant post-processing to extract the output in desired format. A number of prompts were curated for all NLP tasks following the recommended format and instruction pattern proposed by LAraBench (Abdelali et al., 2024). After obtaining a reasonable prompt, we used OpenAI API for GPT 3.5. For Bloomz we ran the model on Google Colab utilizing 16GB GPU and for Llama 2, we used on premises hosted versions utilizing 2X40GB A100 GPUs. Results were post-processed in all cases to align with the test set’s output. the following section elaborates the LLMs (including prompting and post-processing details), NLP Task, Datasets, SOTA Models and evaluation metrics, used in the study.
@@ 2.1 Models
2.1.3 Llama 2
Llama 2 (Touvron et al., 2023) released by Meta is trained on 2 trillion tokens. The model is trained on 89.70% of English content. For evaluation, Llama 2-7b was used due to its open-source availability, and potential for transfer learning and generalization to languages with limited data.
@@ 2.2 Tasks and Datasets
This study has focused on a comprehensive evaluation of pre-trained open and closed LLMs on Urdu NLP tasks. This study utilizes 15 publicly available datasets ( see Table 1) to evaluate 14 Urdu NLP tasks as discussed in the following sections.
2.2.1 Named Entity Recognition
Named Entity Recognition (NER) is a sequence tagging task that involves identifying entities, such as names of people, organizations, locations, dates, etc. For its evaluation, we used the MK-PUCIT dataset and its SOTA model reported in (Kanwal et al., 2019).
Task Dataset Metric GPT 3.5 Bloomz 3B Bloomz 7B1 Llama 2 SOTA Delta
Name Entity Recognition MK-PUCIT Macro-F1 0.55 0.25 0.27 0.15 0.77 0.22
News Categorization COUNTER Macro-F1 0.87 0.58 0.48 0.13 0.7 -0.17
Intent Detection Urdu Web Queries Dataset (UWQ-22) Macro-F1 0.3 0.22 0.18 0.07 0.90 0.60
Hate Speech Detection ISE-Hate corpus Macro-F1 0.72 0.52 0.53 0.48 0.83 0.10
Hate Speech Detection CLE-Hatespeech dataset Macro-F1 0.67 0.35 0.43 0.51 0.98 0.31
Propaganda Detection ProSOUL Macro-F1 0.31 0.47 0.47 0.44 0.83 0.36
Abusive Language Detection HAOSOC - Task A Macro-F1 0.23 0.51 0.47 0.44 0.88 0.37
Threat Detection HAOSOC - Task B Macro-F1 0.49 0.35 0.2 0.47 0.54 0.05
Cyber bullying Identification (Adeeba et al., 2024) Macro-F1 0.19 0.15 0.1 0.06 0.84 0.65
Fake News Detection (Khan et al., 2023) Macro-F1 0.55 0.52 0.51 0.47 0.93 0.38
Hate Speech Categorization ISE-Hate corpus Macro-F1 0.4 0.28 0.15 0.21 0.83 0.43
Text Summarization CORPURES Average Rouge-2 F1 score 0.54 0.46 0.55 0.59 0.57 -0.02
Sentiment Analysis (Muhammad and Burney, 2023) Macro-F1 0.62 0.35 0.33 0.3 0.88 0.26
Sentiment Analysis Corpus of Aspect-based Sentiment for Urdu Political Data Macro-F1 0.31 0.20 0.21 0.13 0.70 0.24
Multi-label Emotion Classification Overview of EmoThreat (Task A) Macro-F1 0.20 0.17 0.26 0.68 0.42
Emotion Classification Urdu Nastalique Emotions Dataset (UNED) Macro-F1 0.32 0.25 0.21 0.18 0.87 0.55
Machine Translation (Quran) English-Urdu Religious Parallel Corpus BLEU 3.75 1.91 2.36 2.49e-78 13.24 9.49
Machine Translation(Bible) English-Urdu Religious Parallel Corpus BLEU 5.96 2.28 2.47 0.097 13.99 8.03
Table 2: Results from zero-shot experiments over 14 tasks with GPT 3.5 , Bloomz-3B, Bloomz-7B1, Llama 2 and SOTA. Bold text indicates the best score.
@@ 2.2.5 Text Summarization
Text summarization involves extracting the most important sentences from a document to create a condensed version retaining essential information. For the evaluation of this task, we used the CORPURES dataset by (Humayoun and Akhtar, 2022).
@@ 2.3 Zero-Shot Setup
For all LLMs; GPT 3.5, Bloomz and Llama 2, we use zero-shot prompting giving natural language instructions describing the task and specify the expected output. Prompts allow LLMs to learn context and narrows the inference space to produces accurate output as further elaborated in the section 2.5.
@@ 2.4 Inference Settings
The inference experiments with Llama2 were conducted using the GPU infrastructure available at the University of Konstanz, Germany. Two parallel NVIDIA A100-PCIE-40GB GPUs were employed, providing a combined computational capacity of 80GB. During the inference of Llama2, nearly 90 percent of the total GPU capacity was utilized. For experiments of GPT-3.5, its original OpenAI API was utilized. The pricing scheme for this API (api, ) involves a cost of \$0.50 per 1 million tokens for input and \$1.50 per 1 million tokens for output. Inference experiments with GPT-3.5 were conducted using Google Colab. Inference experiments with Bloomz’s 3B and 7.1B models, available on huggingface, were also conducted using Google Colab.
@@ 2.5 Prompt Engineering and Post Processing
In our experimentation with different LLMs, we tweaked the prompts based on the models input. Prompts for tasks such as News categorization A.2 and Hate speech Categorization A.11 were challenging because they required outputs from pre-defined ground-truth categories. Prompts for Machine Translation task A.17 had to be engineered so that the model’s output only includes the translated text. Thus optimal prompts were curated by testing against each model on few samples, while ensuring no bias in decision-making.
Despite careful prompting, model responses required post-processing to align with desired outcomes e.g. capitalization ("fake" vs. "Fake"), standardizing output formats ("1. Propaganda" to "1"), and omitting "explanations" and "note" produced with the models’ responses, specifically in Hate speech detection A.5 task. Some model outputs didn’t match desired outcomes, e.g. News categorization included 5 domains i.e. sports, showbiz, foreign , national , business however the models might output out of context domains such as "politics" and "entertainment". Among all the models, Llama 2 required the most output post-processing.
@@ 2.6 SOTA Models
In this study, we benchmark the capabilities of LLMs in a zero-shot scenario by comparing them with SOTA models that employ diverse architectures. These architectures include Capsule NN, Support Vector Machine (SVM), Random Forest (RF), Decision Tree (J48), Sequential Minimal Optimization (SMO), Convolutional Neural Networks (1D-CNN), LSTM with CNN features , Naive Bayes classifier and various multilingual transformer models such as m-BERT and frameworks like XGboost and LGBM.
@@ 2.7 Evaluation Metrics
The evaluation metrics used for the experiments have been kept identical to the one used in the respective state of the art references. They are Macro-F1, Rouge 2 F1 score and BLEU  [^1] https://www.nltk.org/api/nltk.translate.bleu. We have also computed the delta to highlight the differential between best performing LLM’s output with the SOTA model.
@@ 3 Results and Discussion
The results of our experimentation have been summarized in Figure 1. Our results show that LLMs differ in their applicability to different data regimes and tasks. GPT 3.5 surpassed the SOTA model for news categorization in Urdu and Llama 2 for Text Summarizartion. In all other experiments, LLMs remained lower than the SOTA models (reference Table 2). The delta between GPT 3.5 and SOTA model was quit low i.e 0.056 for threat detection task. In comparison with other the open LLMs, GPT 3.5 performed better in majority of the NLP tasks which might be due to its extensive architecture and advanced training techniques, enabling it to effectively generalize across languages and tasks.
Bloomz 3b and Bloomz-7B1 and Llama 2 outperformed GPT 3.5 in four NLP tasks, i.e. Hate speech detection (for CLE HateSpeech dataset), Propaganda Detection, Abusive Language Detection, and Text summarization. One reason for their superior performance on Ethics and NLP tasks compared to GPT 3.5 could be due to their inherent focus on truthfulness, bias and toxicity in the pre-trained models (Touvron et al., 2023). Bloomz had an overall higher Macro F1 score than Llama 2 in majority of evaluation tasks.
Llama 2 mostly demonstrated lower performance in all Urdu NLP tasks when compared to Bloomz except in Text Summarization, threat detection and heat speech detection. This result can be ascribed to the scarcity of non-English data in its pre-training, even though the selected model had twice as many parameters as the Bloomz 3b model.
@@
Figure 1: Average Performance of models as compared to SOTA.
Thus, performance of LLMs significantly depends on well-curated prompts and intelligent post-processing of the outputs. While Llama 2 and Bloomz show a notable performance deficit compared to the SOTA, GPT 3.5 succeeds in mitigating this gap to a certain extent.

### v2

# Benchmarking the Performance of Pre-trained LLMs across Urdu NLP Tasks

@@ Abstract
Large Language Models (LLMs) pre-trained on multilingual data have revolutionized natural language processing research, by transitioning from languages and task specific model pipelines to a single model adapted on a variety of tasks. However majority of existing multilingual NLP benchmarks for LLMs provide evaluation data in only few languages with little linguistic diversity. In addition these benchmarks lack quality assessment against the respective state-of the art models. This study presents an in-depth examination of 7 prominent LLMs: GPT-3.5-turbo, Llama 2-7B-Chat, Llama 3.1-8B, Bloomz 3B, Bloomz 7B1, Ministral-8B and Whisper (Large, medium and small variant) across 17 tasks using 22 datasets, 13.8 hours of speech, in a zero-shot setting, and their performance against state-of-the-art (SOTA) models, has been compared and analyzed. Our experiments show that SOTA models currently outperform encoder-decoder models in majority of Urdu NLP tasks under zero-shot settings. However, comparing Llama 3.1-8B over prior version Llama 2-7B-Chat, we can deduce that with improved language coverage, LLMs can surpass these SOTA models. Our results emphasize that models with fewer parameters but richer language-specific data, like Llama 3.1-8B, often outperform larger models with lower language diversity, such as GPT-3.5, in several tasks.
@@ 1 Introduction
Our study, focuses on evaluating the potential of both closed and open LLMs for supporting Urdu, a low resource language with limited data coverage in LLM’s pre-training. In our experiments we utilize GPT3.5 turbo by OpenAI, Llama 2 and Llama 3.1 by Meta , Bloomz 3B and 7B1 by Big Science, Ministral 8B by Mistral AI and Whisper by OpenAI in zero-shot setting, and perform evaluation on 17 Urdu NLP tasks analyzing their performances with the existing SOTA models. To the best of our knowledge, this is the first in depth evaluation of prominent LLMs in Urdu Language context.
@@
Cyber Bullying Identification Cyberbullying corpus (Adeeba et al., 2024) 12759 2480
@@
Machine Translation(Quran) English-Urdu Religious Parallel Corpus (Jawaid and Zeman, 2011) 6414 200
Machine Translation(Bible) English-Urdu Religious Parallel Corpus (Jawaid and Zeman, 2011) 7957 257
Abstractive Summarization CLE Meeting Corpus (Sadia et al., 2024) 240 10
POS Tagging Sense Tagged CLE Urdu Digest Corpus (Urooj et al., 2014) 100000 22522
ASR (Read Speech) Urdu Speech Corpus (Farooq et al., 2019) - 9.5 hours
ASR (Broadcast) Urdu Broadcast (BC) Corpus(Khan et al., 2021) - 4.3 hours
@@ 2 Approach
For benchmarking of Urdu NLP tasks, we perform experiments using GPT 3.5, Bloomz 3B and Bloomz 7B1 , Llama 2 and Llama 3.1, Ministral 8B and Whisper in zero-shot setting and comparatively analyse the results with the respective SOTA models. Model selection was based on factors like accessibility (open/closed), infrastructure requirement, performance and language support. GPT 3.5 was selected because of its superior performance on English tasks. Among open models, popular multilingual models i.e. Llama 2 , Llama 3.1, Ministral 8B and Bloomz were evaluated for text processing tasks and Whisper models were evaluated for speech recognition task. Due to budget limitations and lack of Urdu data in the pre-training, other closed LLMs models were not investigated.
The evaluation of LLMs involved prompting and significant post-processing to extract the output in desired format. A number of prompts were curated for all NLP tasks following the recommended format and instruction pattern proposed by LAraBench (Abdelali et al., 2024). The prompts for each model were optimized after testing them on a few samples for each task. These prompts have been reported in Appendix. A After obtaining a reasonable prompt, we used the LLM models in different settings. OpenAI’s API was used for GPT 3.5. For Bloomz, we ran the model on Google Colab utilizing 16GB GPU and for Llama 2, Llama 3.1 and Ministral 8B, we used on premises hosted versions utilizing 2X40GB A100 GPUs. Results were post-processed in all cases to align with the test set’s output. The following section elaborates the LLMs (including prompting and post-processing details), NLP Tasks, Datasets, SOTA Models and evaluation metrics, used in the study.
@@ 2.1 Models
2.1.3 Llama 2 and Llama 3.1
Llama 2 (Touvron et al., 2023), released by Meta, is trained on 2 trillion tokens, with 89.70% of its content in English. Llama 3.1 (AI@Meta, 2024), available in three variants with 8 billion, 70 billion and 405 billion parameters, is trained on over 15 trillion tokens. Both models support 8k context lengths. For evaluation, the Llama 2-7b and Llama 3.1-8b models were used due to their open-source availability and potential for transfer learning and generalization to languages with limited data.
2.1.4 Ministral 8B
The Ministral 8B (Mistral AI Team, 2024) is trained on a mixture of multilingual and code datasets, supporting a context window of up to 128k facilitated by an interleaved sliding-window attention mechanism and a vocabulary of 131k. We benchmarked this model due to its open-source availability and its capability for low-memory inference, and its ability to be fine-tuned and adapted to a variety of tasks.
2.1.5 Whisper
Whisper (Radford et al., 2022), an Automatic Speech Recognition (ASR) model developed by OpenAI, is trained on an extensive dataset comprising 680,000 hours of multilingual and multitask supervised data collected from the web. Among the diverse languages included, Whisper incorporates only 104 hours of Urdu speech corpus. For inference , we utilized the small, medium and large variants of the pre-trained Whisper model. The small variant has 12 layers, 12 attention heads, a width of 768 with 244 million parameters. The medium variant is characterized by 24 layers, 16 attention heads, a width of 1024, and consists of 769 million parameters while, the large variant features 32 layers, 20 attention heads, a width of 1280, and comprises 1550 million parameters.
@@ 2.2 Tasks and Datasets
This study has focused on a comprehensive evaluation of pre-trained open and closed LLMs on Urdu NLP tasks. This study utilizes 22 publicly available datasets ( see Table 1) to evaluate 17 Urdu NLP tasks as discussed in the following sections.
2.2.1 Name Entity Recognition
Name Entity Recognition (NER) is a sequence tagging task that involves identifying entities, such as names of people, organizations, locations, dates, etc. For its evaluation, we used the MK-PUCIT dataset and its SOTA model reported in (Kanwal et al., 2019).
Task Dataset Metric GPT 3.5 Bloomz 3B Bloomz 7B1 Llama 2 Llama 3.1 Ministral 8B SOTA Delta
Name Entity Recognition MK-PUCIT Macro-F1 0.55 0.25 0.27 0.15 0.41 0.25 0.77 0.22
News Categorization COUNTER Macro-F1 0.87 0.58 0.48 0.13 0.64 0.67 0.70 -0.17
Intent Detection Urdu Web Queries Dataset (UWQ-22) Macro-F1 0.30 0.22 0.18 0.07 0.42 0.34 0.90 0.56
Hate Speech Detection ISE-Hate corpus Macro-F1 0.72 0.52 0.53 0.48 0.70 0.53 0.83 0.11
Hate Speech Detection CLE-Hatespeech dataset Macro-F1 0.67 0.35 0.43 0.51 0.72 0.54 0.98 0.26
Propaganda Detection ProSOUL Macro-F1 0.31 0.47 0.47 0.44 0.66 0.53 0.83 0.17
Abusive Language Detection HAOSOC - Task A Macro-F1 0.23 0.51 0.47 0.44 0.50 0.48 0.88 0.37
Threat Detection HAOSOC - Task B Macro-F1 0.49 0.35 0.20 0.21 0.40 0.46 0.54 0.05
Cyber Bullying Identification (Adeeba et al., 2024) Macro-F1 0.19 0.15 0.10 0.06 0.22 0.08 0.84 0.41
Fake News Detection (Khan et al., 2023) Macro-F1 0.55 0.52 0.51 0.47 0.72 0.57 0.93 0.21
Hate Speech Categorization ISE-Hate corpus Macro-F1 0.40 0.28 0.15 0.21 0.30 0.22 0.83 0.43
Extractive Summarization CORPURES Average Rouge-2 F1 score 0.54 0.46 0.55 0.59 0.62 0.52 0.57 -0.04
Abstractive Summarization CLE Meeting Corpus Rouge-1 Score (Avg) 0.22 0.02 0.07 0.006 0.24 0.06 0.31 0.07
Sentiment Analysis (Muhammad and Burney, 2023) Macro-F1 0.62 0.35 0.33 0.3 0.44 0.36 0.88 0.26
Sentiment Analysis Corpus of Aspect-based Sentiment for Urdu Political Data Macro-F1 0.31 0.20 0.21 0.13 0.37 0.28 0.70 0.37
Multi-label Emotion Classification Overview of EmoThreat (Task A) Macro-F1 0.20 0.17 0.26 0.40 0.29 0.68 0.28
Emotion Classification Urdu Nastalique Emotions Dataset (UNED) Macro-F1 0.32 0.25 0.21 0.18 0.24 0.41 0.87 0.46
Machine Translation (Quran) English-Urdu Religious Parallel Corpus BLEU 3.75 1.91 2.36 2.49e-78 3.44 0.004 13.24 9.49
Machine Translation(Bible) English-Urdu Religious Parallel Corpus BLEU 5.96 2.28 2.47 0.097 6.43 1.31e-78 13.99 8.03
POS Tagging CLE Urdu POS Tagset Accuracy 0.49 0.11 0.06 0.09 0.31 0.14 0.96 0.47
Table 2: Results from zero-shot experiments of GPT 3.5, Bloomz 3B, Bloomz 7B1, Llama 2, Llama 3.1 and Ministral 8B Models Compared to SOTA over NLP tasks. Bold text indicates the best score among models.
@@ 2.2.5 Text Summarization
Text summarization involves extracting the most important sentences from a document to create a condensed version retaining essential information. We evaluated the LLMs on:
  • Extractive Summarization
    Extractive summarization condenses text by selecting and combining key sentences directly from the original content. For the evaluation of this task, we used the CORPURES dataset by (Humayoun and Akhtar, 2022).
  • Abstractive Summarization
    Abstractive summarization generates concise summaries by understanding and paraphrasing the core meaning of a text into new, shorter sentences. For its evaluation, we have used CLE Meeting Corpus and its SOTA available in (Sadia et al., 2024).
@@ 2.2.7 Machine Translation
2.2.8 Part of Speech (POS) Tagging
POS tagging is a fundamental task in NLP that involves labeling each word in a sentence with its corresponding part of speech, such as noun, verb, adjective, etc. To evaluate this task we have used the CLE Urdu POS Tagset with the SOTA reported in (Ahmed et al., 2014).
2.2.9 Automatic Speech Recognition (ASR)
ASR automatically converts spoken language into text. For its evaluation, we utilized the small, medium and large variant of the pre-trained Whisper model (Radford et al., 2022). We benchmarked this model against the SOTA models using its pre-trained weights for both broadcast and read speech recognition tasks using following two corpora:
  • Urdu Broadcast (BC) corpus: A broadcast speech corpus (Khan et al., 2021) with 4.3 hours of data from 25 speakers (14 males and 11 females). This dataset includes recordings from five different broadcast channels and YouTube, covering genres such as entertainment, health and science, current affairs, and politics.
  • Urdu Speech corpus: A read speech corpus (Farooq et al., 2019) consisting of 9.5 hours of Urdu speech from 62 speakers. The dataset is balanced in terms of gender and recording channels.
Task Domain Metric Whisper (Large) Whisper (Medium) Whisper (Small) SOTA Delta
ASR Read Speech WER 23.51 27.88 36.90 16.94 -6.57
ASR Broadcast WER 27.97 35.57 42.57 18.59 -9.38
Table 3: Performance Metrix of ASR for Whisper Large, Medium and Small Models Compared to SOTA.
@@ 2.3 Zero-Shot Setup
For all LLMs; GPT 3.5, Bloomz 3b and 7b, Llama 2 and Llama 3.1 and Ministral 8B we use zero-shot prompting giving natural language instructions describing the task and specify the expected output. Prompts allow LLMs to learn context and narrows the inference space to produces accurate output as further elaborated in the section 2.5.
@@ 2.4 Inference Settings
The inference experiments for Llama 2, Llama 3.1 and Ministral 8B were conducted using two parallel NVIDIA A100-PCIE-40GB GPUs, providing a combined computational capacity of 80GB. During the inference, nearly 90 percent of the total GPU capacity was utilized. For experiments of GPT-3.5, API from OpenAI was utilized. Inference experiments with GPT-3.5 were conducted using Google Colab. Inference experiments with Bloomz’s 3B and 7.1B models, available on huggingface, were also conducted using Google Colab. For Speech processing experiments using Whisper, two NVIDIA RTX3060-12GB GPUs were employed, providing a combined computational capacity of 24GB.
@@ 2.5 Prompt Engineering and Post Processing
In our experimentation with different LLMs, we tweaked the prompts based on the models input. Prompts for tasks such as News categorization A.2 and Hate speech Categorization A.11 were challenging because they required outputs from pre-defined ground-truth categories. Prompts for Machine Translation task A.19 had to be engineered so that the model’s output only includes the translated text. Thus optimal prompts were curated by testing against each model on few samples, while ensuring no bias in decision-making.
Despite careful prompting, model responses required post-processing to align with desired outcomes e.g. capitalization ("fake" vs. "Fake"), standardizing output formats ("1. Propaganda" to "1"), and omitting "explanations" and "note" produced with the models’ responses, specifically in Hate speech detection A.5 task. Some model outputs didn’t match desired outcomes, e.g. News categorization included 5 domains i.e. sports, showbiz, foreign , national , business however the models output out of context domains such as "politics" and "entertainment". Among all the models, Llama 2 required the most output post-processing.
@@ 2.6 SOTA Models
In this study, we benchmark the capabilities of LLMs in a zero-shot scenario by comparing them with SOTA models as reported in respective studies. These SOTA models employed diverse architectures including Capsule NN, Support Vector Machine (SVM), Random Forest (RF), Decision Tree (J48), Sequential Minimal Optimization (SMO), Convolutional Neural Networks (1D-CNN), LSTM with CNN features , Naive Bayes classifier and various multilingual transformer models such as m-BERT and frameworks like XGboost and LGBM.
@@ 2.7 Evaluation Metrics
The evaluation metrics used for the experiments have been kept identical to the one used in the respective state of the art references. They are Macro-F1, Rouge 2 F1 score, BLEU  [^1] https://www.nltk.org/api/nltk.translate.bleu, accuracy and Word Error Rate (WER). We have also computed the delta to highlight the differential between best performing LLM’s output with the SOTA model.
@@ 3 Results and Discussion
The results on text processing tasks of our experimentation have been summarized in Figure 1. The Figure presents a grid of bar graphs for each NLP task, with the y-axis showing evaluation metrics specific to each task. For classification and detection tasks, the y-axis represents the macro F1 score. For summarization tasks, it shows the average ROUGE-2 score, while for machine translation tasks, it displays the BLEU score. Each model is represented by a distinct color bar , as indicated in the Figure’s legend, which is kept consistent across all tasks, with the bar of SOTA providing a reference point for comparison. Missing bars in certain tasks indicate that the model outputs were effectively zero (e.g., in Table 2 value is 2.49e-78 for Llama 2 on the Machine Translation (Quran) task, and 1.31e-78 for Mistral 8B on the Machine Translation (Bible) task), reflecting negligible performance.
@@
Our results show that LLMs differ in their applicability to different data regimes and tasks. LLM models were able to surpassed the SOTA model for news categorization with GPT 3.5 and Llama 3.1 for Extractive Summarization. In all other experiments, LLMs remained lower than the SOTA models (reference Table 2). Across all experiments, Llama 3.1 outperformed in 10 of the 17 tasks, while GPT-3.5 excelled in 8 tasks. In comparison, Bloomz and Ministral 8B each led in only one task. The minimum delta obtained was 0.05 between GPT 3.5 and SOTA model for threat detection task. In comparison with other the open LLMs, Llama 3.1 performed better in majority of the NLP tasks which is due to its extensive multilingual data, architecture and advanced training techniques, enabling it to effectively generalize across languages and tasks.
Figure 1: The performance of different models in zero-shot scenario as compared to SOTA. Missing bars in some tasks mean that the specific model cannot perform the specified task.
In choice and evaluation of LLMs, Bloomz 3B and Bloomz 7B1 were initially chosen for experimentation due to their early introduction and multilingual capabilities. However, they have not kept pace with advancements seen in other models like Llama. Analysis reveals that there is no significant performance efficiency gained from transitioning from Bloomz-3B to Bloomz-7B1 as evident from Table 2. In contrast, the performance of Llama models has notably improved, particularly from Llama 2 to Llama 3.1, indicating a more effective evolution in their design and capabilities.
Based on our evaluations, the top two performing models for NLP tasks are GPT-3.5 and Llama 3.1 with comparable performances as evident from Figure 1. Llama 3.1 outperformed GPT 3.5 in 11 NLP tasks and was even better than SOTA in Extractive Summarization. On the other hand, GPT 3.5 was better than Llama 3.1 in 8 tasks and surpassed SOTA in News Categorization task. Llama 3.1’s superior performance is due to the increased coverage of Non-English data in the model as well as increased amount of pretraining data i.e. 15 trillion tokens.
The performance of Ministral 8B is comparable to GPT 3.5 and Llama 3.1. It outperformed GPT 3.5 in 6 NLP Tasks i.e. Intent Detection, Propaganda Detection, Abusive Language Detection, Fake News Detection, Multi Label Emotion Classification and Emotion Classification. And outperformed Llama 3.1 in News Categorization and it was best among all models in Emotion Classification. However its performance was quite inadequate on generation tasks like Machine Translation (reference Figure 1). Overall its performance is good in detection tasks and is attributed to its interleaved sliding-window attention mechanism, which enables it to efficiently handle extended contexts with reduced memory usage. Detection tasks such as Fake News Detection and Propaganda Detection often require recognizing patterns across longer texts. This enhanced ability to retain and utilize extended contextual information allowed Ministral 8B to excel in detection tasks.
The results of Speech Processing tasks are summarized in Table 3. The analysis of the results indicates that the SOTA models, which were trained on a larger corpus of Urdu data, outperformed all variants of the Whisper model across the evaluated datasets. These findings suggest that while the SOTA models trained on more extensive Urdu datasets exhibited superior performance, the larger variant of the Whisper also demonstrated improved performance compared to its medium and small counterpart, underscoring the importance of model size and complexity in ASR tasks. The negative delta values indicate that the model’s performance falls below the SOTA benchmarks, highlighting a gap to be addressed.
@@
Thus, performance of LLMs significantly depends on well-curated prompts and intelligent post-processing of the outputs. While Llama 2 and Bloomz show a notable performance deficit compared to the SOTA, the newer Llama version i.e. Llama 3.1 and GPT 3.5 succeeds in mitigating this gap to a considerable extent.